40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2186     Entry time: Tue May 15 14:29:47 2018
Author: awade 
Type: DailyProgress 
Category: TempCtrl 
Subject: Beat drift over the last few weeks 

We haven't been paying attention to the beat frequency or its control in the last few weeks.  Yesterday it was apparent that the beat had drifted to ~2 GHz.  When I tuned the differential heating to ~0.42 W with a common offset of 0.56831 W I found that this had almost no effect on the beat note.

Initially I thought maybe the shield heater driver box was broken.  On closer inspection it seems that the Acromag IOC channels had been moved to a different output channel and the IOC process was only very recently reset to enact those changes.  

I've relabeled the cables and BNC ports to make it clear which components should plug into which.  I've set the differential heating to 0.422991 W and both lasers seem to be tracking in the right direction towards a 26 MHz beat note now.  It will take a few more hours to get close enough to stabilized. Currently at 250 MHz.

ELOG V3.1.3-